# SIGMUS: Semantic Integration for Knowledge Graphs in Multimodal Urban Spaces**footnotemark: *

### v1

@@ 1. Introduction
footnotetext: The author holds concurrent appointments as Amazon Scholar and Professor at UCLA, but the work in this paper is not associated with Amazon.
@@
Recently, Large Language Models (LLMs) have been shown to demonstrate significant world knowledge about sensory data (xu_penetrative_2024, ; yang_drhouse_2024, ). We hypothesize that LLMs, when used to reason about incidents and sensory data, provide valuable insights into the emergence of real-world incidents and their detection through sensory data. Importantly, the ability to generate meaningful relationships between incidents and different sensory sources (e.g. CCTV, traffic measurements, news articles) requires a significant degree of world knowledge. In this work, we show the effectiveness of LLMs at the qualitative reasoning required to organize and create relationships between incidents and sensory data.
@@ 2.1. Multimodal Knowledge Graphs
Existing works on multimodal knowledge graphs typically involve both semi-structured text and complex unstructured data such as images. (kannan_multimodal_2020, ) focuses on extracting text, images, and source code from academic publications and creating their respective graphs. They apply a separate processing pipeline for each modality and attempt to align different entities into the Computer Science Ontology (CSO) (noauthor_cso_nodate, ). (wang_tiva-kg_2023, ) takes an existing knowledge graph (ConceptNet (speer2017conceptnet, )) and annotates it with additional attributes for images, sound, and video. (usmani_towards_2023, ) describes an ontology for structuring knowledge collected by sensors in smart cities. However, it lacks a concrete implementation and doesn’t collect event information necessary for smart city applications.
@@ 2.2. Multimodal Data Integration
The concept of unifying data sources which are diverse in type into a coherent view for monitoring or prediction has also been explored. In the web domain, a sensor web (broring2011new, ) describes an infrastructure that enables discovery and integration of sensors and their data. An example of such a system is Senseweb (grosky2007senseweb, ), which is a web platform that ingests data from different sensory data streams and arranges them into a map view. Some tools, such as Graph of Things (le2016graph, ) describe an idea similar to ours, where they aim to construct a live knowledge graph of various IoT and social media data, and creating connections between heterogenous data sources using several ontologies. While these platforms and tools share a similar goal to ours, they lack the ability to produce inferences and correlations between information which has historically required human intervention to directly process such correlations or create flexible pipelines to represent such reasoning. It’s also worth mentioning that while multimodal data fusion also brings together data from different sources, its objective is often task-specific and typically doesn’t use human knowledge as explicitly as data integration might (such as the use of an ontology or linking against an external knowledge graph). Thus we won’t be including related work from this field.
@@ 2.3. LLMs for Knowledge Graphs
(holur_creating_2024, ) iteratively creates a semantic map of events occurring across news documents, and uses the LLM to fill out actors, relationships, and states given the raw text document. (giglou_llms4ol_2023, ) attempts to create an ontology from a document of terms, using some internal knowledge of the world to decide taxonomic and non-taxonomic relations. (zhu_llms_2024, ) provides knowledge graphs as answers to users questions by searching for relevant information and organizing it into facts represented as triples.
@@
(xu_generate–graph_2024, ) uses an LLM in a Retrieval-Augmented-Generation (RAG) style to query a knowledge graph and answer a user query. In particular, it converts a natural language question into the appropriate query language, and passes the response as part of the context into the LLM for question answering. (toro_dynamic_2024, ) uses additional text from online sources to fill out empty fields in some given knowledge templates. (zaitoun_can_2023, ) uses an existing ontology and seeks to extend it with new relationships between entities given a collection of raw documents.
(adam_traceable_2024, ) uses ground truth documents in order to verify triples found in knowledge graphs using LLMs. (boylan_kgvalidator_2024, ) works with several additional sources of obtaining evidence: web searches, LLM internal knowledge, and reference knowledge graphs.
@@
(edge_local_2025, ) stores nodes from a knowledge graph as document embeddings which are used at query time for RAG. (deng_graphvis_nodate, ) is given a graph as an input, either as a textual description or an image, along with a user query. (wu_retrieve-rewrite-answer_2023, ) aims to improve Q/A performance by rewriting extracted triples from a KG into more descriptive, readable text. (sui_can_2025, ) uses both a knowledge graph as well as some degree of the LLM’s internal reasoning to answer more open ended Q/A questions. (feng_knowledge_2023, ; sun_think_graph_2024, ; jiang_structgpt_2023, ; jin_graph_2024, ) allows for querying the knowledge graph by iteratively traversing the KG. These works treat the LLM as a planner for obtaining the necessary information to answer a question.
@@ 2.4. Ontologies for Multimodal Sensory Knowledge Graphs
Several ontologies were found to be relevant to our scenarios, and are particularly useful when discussing Semantic Web related platforms. The Brick schema (noauthor_brick_nodate, ) describes an RDF-written ontology, designed to represent resources in building management systems. It includes concepts covering individual sensory devices as well as locations within buildings, represented by logical relationships. The Smart Applications Reference Ontology (SAREF) (noauthor_saref_nodate, ) and its wearable version (SAREF4WEAR) (noauthor_saref4wear_nodate, ) both describe sensor-specific characteristics, including properties of the device, measurements, and real world features to be measured. The Sensor, Observation, Sample, Actuator Ontology (SOSA) (noauthor_semantic_nodate, ) provides a similar sensor-centric view, describing features of interest, properties of sensors, as well as measurements.
@@
While there exists some works on smart city modeling, such as (noauthor_azureopendigitaltwins-smartcities_2025, ; espinoza-arias_ontological_2019, ; noauthor_smart_nodate, ), these works either do not involve the same notion of events that we wish to model, lack sensor-specificity, or target simpler structured sensor readings and not more complex information such as actors within news articles.
@@ 3.1. Terminology
Table 1 describes the different terms we use when organizing data in urban spaces. The notion of incidents is motivated by the need for larger occurrences influencing cities or states (such as natural disasters, major entertainment events, etc) to be represented in our knowledge graph, allowing us to identify more specific components contributing to them. Reports are an abstraction where sensory data (and its associated metadata like time, accuracy, and location), as well as human-generated knowledge (such as news and tweets) can be integrated in our knowledge graph. In our system, a report is an individual news article, image, or sample from a time series. Observers can be used to identify the entity publishing a particular report. For sensors, this can be a specific sensor ID, while for human-generated knowledge this can be an organization name (such as a news station) or username. Modality represents a specific modality of a report. This allows us to separate out different types of data present in a report possessing different properties. For example, a news report may use images, audio, and text. Images yield a caption property describing the image, audio describes the loudness of a scene, and text reveals the named actors involved an event. Aggregators are the platforms from which we obtain data from (such as X.com, Caltrans CCTV, etc), and may vary in how trustworthy its knowledge is depending on the nature of its observers (though currently we do not model these attributes). Events allow us to describe interactions between two parties, such as ”accuse of a crime”, or ”protest” (where interactions are described under CAMEO codebook (noauthor_data_nodate, )). In our current implementation, events are specific to news articles, but other modalities may also reveal specific interactions between parties. Actors are the parties involved with an event, where both individual people and larger organizations of people may be named as actors.
@@
the Federal Emergency Management Agency (FEMA) (noauthor_glossary_nodate, ). In our system, incidents may be linked to
@@
types of cooperative behaviors. We use the Conflict and Mediation Event Observations (CAMEO) codebook (noauthor_data_nodate, ) to describe
@@
(CAMEO) codebook (noauthor_data_nodate, ) to describe our actors. However, while CAMEO is typically focused on larger groups
@@ 3.2. The SIGMUS ontology
  • dcterms: identifier is used as a reference to a particular resource represented as some unique key or string, taken from the Dublin Core Terms ontology (noauthor_dcmi_nodate, ).
  • rdfs: label and rdfs: comment are used to describe a human readable string and a description, respectively. These are taken from the RDF Schema (noauthor_rdf_nodate, ).
  • time properties are taken from the OWL Time ontology (hobbs_ontology_2004, ), allowing us to represent the start and end time of events.
  • sosa and qudt are both part of the Semantic Sensor Network Ontology (noauthor_semantic_nodate, ), which we use to describe the properties and units of sensory measurements. For more complex data (like images and audio), the Value field is treated as a filepath to the data rather than directly encoding it in the knowledge graph instance.
@@
  • The sigmus: cameoEvent property represents the codes from CAMEO (noauthor_data_nodate, ) that can be mapped to specific event terms and descriptions (e.g. assassinations) or actors (e.g. government).
@@ 4.1. Data sources and Ingestion
GDELT events (noauthor_gdelt_nodate, ) news article URLs, named actors, CAMEO event codes, Geo names Zip files of CSVs in 15 minute intervals 15min 15min Mentions of Los Angeles
Caltrans PeMS (california_pems_nodate, ) Traffic incidents, vehicle speeds, vehicle occupancy Zip files of CSVs updated daily 1 day 1 day incidents and sensors in LA
Caltrans CCTV (noauthor_caltrans_nodate, ) images from highway cameras Live video streams 15min None Cameras in LA county
ALERT California (noauthor_alertcalifornia_nodate, ) images from wildfire cameras Live video streams 30min None Cameras in LA county
Noise Planet (noauthor_noise-planet_nodate, ) noise levels in db(A) Zip files updated daily 1 day 1 day Measurements from LA
SCEDC Seismic (noauthor_southern_nodate, ) seismic waveform data from earthquake monitoring stations Zip files of waveform data updated daily 1 day 1 day Seismic stations in LA county
OpenWeather (noauthor_current_nodate, ) wind speed, precipitation, and weather descriptions API-based access, updated ẽvery 10 minutes 1 hour None Weather stations in LA county
PurpleAir (noauthor_real-time_nodate, ) PM2.5 API-based access, updated ẽvery 2 minutes 1 hour None Outdoor sensors in LA county
@@
Citizen App (noauthor_citizen_nodate, ) Text alerts, alert category, location of alert App alerts (event-based) 15min 15min Alerts about locations in LA county
@@ 4.2. Knowledge Graph and Integration
Actor and Event Parsing uses LLMs to identify any real world organizations and persons involved in the text, as well as classifying these into the corresponding CAMEO term (noauthor_data_nodate, ). In addition, it must categorize the event occurring between the actors using CAMEO event terms (noauthor_data_nodate, ), as well as identifying any alerts or emergencies using the Common Alerting Protocol (CAP) (noauthor_common_nodate, ), if any. An high level example of this process is shown in Figure 4.
@@ 5. Evaluation
We evaluate SIGMUS using two LLMs. One LLM is a DeepSeek-R1 70B model, which we operate on an H100 GPU server hosted on our local premises. Another is the cloud-based OpenAI GPT-4o model. The reason for these two different models is mainly due to quality of the output, with the GPT-4o performing significantly better than the DeepSeek model on the cross-modailty linking task. For that reason, we rely on the GPT-4o model for the cross-modality linking task, while for all other tasks we utilize DeepSeek-R1. For Visual Event classification, we rely on a Visual Large Language Model called NVILA (liu_nvila_2025, ), which also runs on the local H100 server. During the Incident Merging and Organization task, we utilize a vector database called Marqo (noauthor_marqo_nodate, ), which allows us to find similar incidents based on their textual embeddings. Our data warehouse uses TimescaleDB for its time series storage, and SeaweedFS for the blob storage. All ingested data is organized and visualized in Neo4j, the knowledge graph used in SIGMUS. In our evaluation, data storage services (i.e. time series, blob, vector, KG) operate on the same desktop PC.
@@
As mentioned in section 3, a single report is treated as an individual news article, image, or row in a table. Note that depending on the modality, the processing and knowledge graph latency may be affected. For example, GDELT events (being a text modality) involves calling an LLM to identify actors, events, and incidents (which makes the bulk of the processing latency). However, when inserting such data into a knowledge graph, we have to disambiguate names which refer to the same entity. These are two kinds of disambiguation processes which occur when we insert items into the knowledge graph: Actor disambiguation and Incident disambiguation. Note that because event names are canonical (generated from (noauthor_data_nodate, )), we do not perform disambiguation for these.
@@
Incident Disambiguation When inserting incidents into the knowledge graph, sysname identifies the top similar existing incidents based on both its label (e.g. ”2025 Pacific Palisades Wildfire”) as well as the original text from the article. This process of ranking the top similar incidents is done via vector search in Marqo (noauthor_marqo_nodate, ), which produces embeddings of each incident and allows for similarity searching given a query. After obtaining the top-k similar incidents, our system prompts the LLM to identify if the inserted incident already exists in the top-k. In addition, the LLM also identifies if the inserted incident related to another incident, and we mainly focus on ”IS_PART_OF” as a possible relationship between incidents. This process of inserting incidents takes 7.8677 seconds on average.

### v2

@@ 1. Introduction
Recently, Large Language Models (LLMs) have been shown to demonstrate significant world knowledge about sensory data (Xu et al., 2024a; Yang et al., 2024). We hypothesize that LLMs, when used to reason about incidents and sensory data, provide valuable insights into the emergence of real-world incidents and their detection through sensory data. Importantly, the ability to generate meaningful relationships between incidents and different sensory sources (e.g. CCTV, traffic measurements, news articles) requires a significant degree of world knowledge. In this work, we show the effectiveness of LLMs at the qualitative reasoning required to organize and create relationships between incidents and sensory data.
@@ 2.1. Multimodal Knowledge Graphs
Existing works on multimodal knowledge graphs typically involve both semi-structured text and complex unstructured data such as images. (Kannan et al., 2020) focuses on extracting text, images, and source code from academic publications and creating their respective graphs. They apply a separate processing pipeline for each modality and attempt to align different entities into the Computer Science Ontology (CSO) (10). (Wang et al., 2023) takes an existing knowledge graph (ConceptNet (Speer et al., 2017)) and annotates it with additional attributes for images, sound, and video. (Usmani et al., 2023) describes an ontology for structuring knowledge collected by sensors in smart cities. However, it lacks a concrete implementation and doesn’t collect event information necessary for smart city applications.
@@ 2.2. Multimodal Data Integration
The concept of unifying data sources which are diverse in type into a coherent view for monitoring or prediction has also been explored. In the web domain, a sensor web (Bröring et al., 2011) describes an infrastructure that enables discovery and integration of sensors and their data. An example of such a system is Senseweb (Grosky et al., 2007), which is a web platform that ingests data from different sensory data streams and arranges them into a map view. Some tools, such as Graph of Things (Le-Phuoc et al., 2016) describe an idea similar to ours, where they aim to construct a live knowledge graph of various IoT and social media data, and creating connections between heterogenous data sources using several ontologies. While these platforms and tools share a similar goal to ours, they lack the ability to produce inferences and correlations between information which has historically required human intervention to directly process such correlations or create flexible pipelines to represent such reasoning. It’s also worth mentioning that while multimodal data fusion also brings together data from different sources, its objective is often task-specific and typically doesn’t use human knowledge as explicitly as data integration might (such as the use of an ontology or linking against an external knowledge graph). Thus we won’t be including related work from this field.
@@ 2.3. LLMs for Knowledge Graphs
(Holur et al., 2024) iteratively creates a semantic map of events occurring across news documents, and uses the LLM to fill out actors, relationships, and states given the raw text document. (Giglou et al., 2023) attempts to create an ontology from a document of terms, using some internal knowledge of the world to decide taxonomic and non-taxonomic relations. (Zhu et al., 2024) provides knowledge graphs as answers to users questions by searching for relevant information and organizing it into facts represented as triples.
@@
(Xu et al., 2024b) uses an LLM in a Retrieval-Augmented-Generation (RAG) style to query a knowledge graph and answer a user query. In particular, it converts a natural language question into the appropriate query language, and passes the response as part of the context into the LLM for question answering. (Toro et al., 2024) uses additional text from online sources to fill out empty fields in some given knowledge templates. (Zaitoun et al., 2023) uses an existing ontology and seeks to extend it with new relationships between entities given a collection of raw documents.
(Adam and Kliegr, 2024) uses ground truth documents in order to verify triples found in knowledge graphs using LLMs. (Boylan et al., 2024) works with several additional sources of obtaining evidence: web searches, LLM internal knowledge, and reference knowledge graphs.
@@
(Edge et al., 2025) stores nodes from a knowledge graph as document embeddings which are used at query time for RAG. (Deng et al., ) is given a graph as an input, either as a textual description or an image, along with a user query. (Wu et al., 2023) aims to improve Q/A performance by rewriting extracted triples from a KG into more descriptive, readable text. (Sui et al., 2025) uses both a knowledge graph as well as some degree of the LLM’s internal reasoning to answer more open ended Q/A questions. (Feng et al., 2023; Sun et al., 2024; Jiang et al., 2023; Jin et al., 2024) allows for querying the knowledge graph by iteratively traversing the KG. These works treat the LLM as a planner for obtaining the necessary information to answer a question.
@@ 2.4. Ontologies for Multimodal Sensory Knowledge Graphs
Several ontologies were found to be relevant to our scenarios, and are particularly useful when discussing Semantic Web related platforms. The Brick schema (23) describes an RDF-written ontology, designed to represent resources in building management systems. It includes concepts covering individual sensory devices as well as locations within buildings, represented by logical relationships. The Smart Applications Reference Ontology (SAREF) (33) and its wearable version (SAREF4WEAR) (34) both describe sensor-specific characteristics, including properties of the device, measurements, and real world features to be measured. The Sensor, Observation, Sample, Actuator Ontology (SOSA) (35) provides a similar sensor-centric view, describing features of interest, properties of sensors, as well as measurements.
@@
While there exists some works on smart city modeling, such as (3; 16; 36), these works either do not involve the same notion of events that we wish to model, lack sensor-specificity, or target simpler structured sensor readings and not more complex information such as actors within news articles.
@@ 3.1. Terminology
Table 1 describes the different terms we use when organizing data in urban spaces. The notion of incidents is motivated by the need for larger occurrences influencing cities or states (such as natural disasters, major entertainment events, etc) to be represented in our knowledge graph, allowing us to identify more specific components contributing to them. Reports are an abstraction where sensory data (and its associated metadata like time, accuracy, and location), as well as human-generated knowledge (such as news and tweets) can be integrated in our knowledge graph. In our system, a report is an individual news article, image, or sample from a time series. Observers can be used to identify the entity publishing a particular report. For sensors, this can be a specific sensor ID, while for human-generated knowledge this can be an organization name (such as a news station) or username. Modality represents a specific modality of a report. This allows us to separate out different types of data present in a report possessing different properties. For example, a news report may use images, audio, and text. Images yield a caption property describing the image, audio describes the loudness of a scene, and text reveals the named actors involved an event. Aggregators are the platforms from which we obtain data from (such as X.com, Caltrans CCTV, etc), and may vary in how trustworthy its knowledge is depending on the nature of its observers (though currently we do not model these attributes). Events allow us to describe interactions between two parties, such as ”accuse of a crime”, or ”protest” (where interactions are described under CAMEO codebook (12)). In our current implementation, events are specific to news articles, but other modalities may also reveal specific interactions between parties. Actors are the parties involved with an event, where both individual people and larger organizations of people may be named as actors.
@@
the Federal Emergency Management Agency (FEMA) (19). In our system, incidents may be linked to
@@
types of cooperative behaviors. We use the Conflict and Mediation Event Observations (CAMEO) codebook (12) to describe
@@
(CAMEO) codebook (12) to describe our actors. However, while CAMEO is typically focused on larger groups
@@ 3.2. The SIGMUS ontology
  • dcterms: identifier is used as a reference to a particular resource represented as some unique key or string, taken from the Dublin Core Terms ontology (13).
  • rdfs: label and rdfs: comment are used to describe a human readable string and a description, respectively. These are taken from the RDF Schema (31).
  • time properties are taken from the OWL Time ontology (Hobbs and Pan, 2004), allowing us to represent the start and end time of events.
  • sosa and qudt are both part of the Semantic Sensor Network Ontology (35), which we use to describe the properties and units of sensory measurements. For more complex data (like images and audio), the Value field is treated as a filepath to the data rather than directly encoding it in the knowledge graph instance.
@@
  • The sigmus: cameoEvent property represents the codes from CAMEO (12) that can be mapped to specific event terms and descriptions (e.g. assassinations) or actors (e.g. government).
@@ 4.1. Data sources and Ingestion
GDELT events (41) news article URLs, named actors, CAMEO event codes, Geo names Zip files of CSVs in 15 minute intervals 15min 15min Mentions of Los Angeles
Caltrans PeMS (California, ) Traffic incidents, vehicle speeds, vehicle occupancy Zip files of CSVs updated daily 1 day 1 day incidents and sensors in LA
Caltrans CCTV (7) images from highway cameras Live video streams 15min None Cameras in LA county
ALERT California (2) images from wildfire cameras Live video streams 30min None Cameras in LA county
Noise Planet (30) noise levels in db(A) Zip files updated daily 1 day 1 day Measurements from LA
SCEDC Seismic (37) seismic waveform data from earthquake monitoring stations Zip files of waveform data updated daily 1 day 1 day Seismic stations in LA county
OpenWeather (11) wind speed, precipitation, and weather descriptions API-based access, updated ẽvery 10 minutes 1 hour None Weather stations in LA county
PurpleAir (32) PM2.5 API-based access, updated ẽvery 2 minutes 1 hour None Outdoor sensors in LA county
@@
Citizen App (8) Text alerts, alert category, location of alert App alerts (event-based) 15min 15min Alerts about locations in LA county
@@ 4.2. Knowledge Graph and Integration
Actor and Event Parsing uses LLMs to identify any real world organizations and persons involved in the text, as well as classifying these into the corresponding CAMEO term (12). In addition, it must categorize the event occurring between the actors using CAMEO event terms (12), as well as identifying any alerts or emergencies using the Common Alerting Protocol (CAP) (9), if any. An high level example of this process is shown in Figure 4.
@@ 5. Evaluation
We evaluate SIGMUS using two LLMs. One LLM is a DeepSeek-R1 70B model, which we operate on an H100 GPU server hosted on our local premises. Another is the cloud-based OpenAI GPT-4o model. The reason for these two different models is mainly due to quality of the output, with the GPT-4o performing significantly better than the DeepSeek model on the cross-modailty linking task. For that reason, we rely on the GPT-4o model for the cross-modality linking task, while for all other tasks we utilize DeepSeek-R1. For Visual Event classification, we rely on a Visual Large Language Model called NVILA (Liu et al., 2025), which also runs on the local H100 server. During the Incident Merging and Organization task, we utilize a vector database called Marqo (29), which allows us to find similar incidents based on their textual embeddings. Our data warehouse uses TimescaleDB for its time series storage, and SeaweedFS for the blob storage. All ingested data is organized and visualized in Neo4j, the knowledge graph used in SIGMUS. In our evaluation, data storage services (i.e. time series, blob, vector, KG) operate on the same desktop PC.
@@
As mentioned in section 3, a single report is treated as an individual news article, image, or row in a table. Note that depending on the modality, the processing and knowledge graph latency may be affected. For example, GDELT events (being a text modality) involves calling an LLM to identify actors, events, and incidents (which makes the bulk of the processing latency). However, when inserting such data into a knowledge graph, we have to disambiguate names which refer to the same entity. These are two kinds of disambiguation processes which occur when we insert items into the knowledge graph: Actor disambiguation and Incident disambiguation. Note that because event names are canonical (generated from (12)), we do not perform disambiguation for these.
@@
Incident Disambiguation When inserting incidents into the knowledge graph, sysname identifies the top similar existing incidents based on both its label (e.g. ”2025 Pacific Palisades Wildfire”) as well as the original text from the article. This process of ranking the top similar incidents is done via vector search in Marqo (29), which produces embeddings of each incident and allows for similarity searching given a query. After obtaining the top-k similar incidents, our system prompts the LLM to identify if the inserted incident already exists in the top-k. In addition, the LLM also identifies if the inserted incident related to another incident, and we mainly focus on ”IS_PART_OF” as a possible relationship between incidents. This process of inserting incidents takes 7.8677 seconds on average.